5. Solve the given second order differential equation by undetermined coefficients. y – 3y = 8e3x + 4 sin x
Answer #1

Given that (S) y-ay = 86%P+ 4xion 400g -- 3 dy6C) 37 + 4 xime Joca DY) - 3DY (X) = Q3 + 4 siny & ex + 4 sinne (02-30) y® Lofor solving &,,22 423 Compute d Yp(x) d (2, e30 x + dy lego x + Tgkin dre Jx 2, 3 + 34, _ Lexinfr) + <3 Color) and of ople) 22 sinne ... YAO) = 38 m) + scono 6 Con 5 the General Polcation. y%= 40 + yoles 2 37 xt Glostry - 3 Air of 460) = Carex oling

If you have any doubt then please comment.